HUBBLE SPACE TELESCOPE DAILY REPORT #2044 PERIOD COVERED: 0000Z (UTC) 01/21/98 - 0000Z (UTC) 01/22/98 Daily Status Report as of 022/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7711 (STIS CCD G750L Fringing Flats.) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD flat-field fringing in the G750L grating at its nominal wavelength setting {7751A}. WAVECALs were also taken at the same gain and wavelength setting. Typical S/N = 100 for each integration, and gain = 4 on each. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of NIC/1/2/3 7235 (Observations of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Hubble Deep Field area HDF-123645. This program observes a portion of the Hubble Deep Field with all three NICMOS cameras in the F160W and F110W filters to produce infrared images equivalent in depth to the existing WF/PC-2 HDF images. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/MA1/MA2 7921 (STIS MAMA Pure Parallels for NIC/3 HDF - Spec) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make pure parallel, ultraviolet observations to be used in the NIC/3 HDF campaign. The MAMA images will be used to search for Lyman-break galaxies at z>2.0, and z>1.0. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Three Sets of STIS/MA1/MA2 7920 (STIS MAMA Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make parallel field, ultraviolet imaging to search for Lyman-break galaxies at z>2.0, and z>1.0. The observations were completed as planned, and no problems were reported. 1.6 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Four Sets of STIS/CCD 7900 (STIS CCD Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (CCD) was used to make pure parallel, ultraviolet, slitless, observations with the NIC/3 Camera. This proposal will run during the orbits devoted to NICMOS HDF prime proposal 7235. The observations were completed as scheduled, and no anomalies were reported. 1.8 Completed Three Sets of NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/CCD 7918 (STIS CCD Pure Parallels for NIC/3 HDF Spec) The Space Telescope Imaging Spectrograph (CCD) was used to make pure parallel, ultraviolet, slitless, spectrographic observations with the prime Hubble Deep field observations. This proposal will run during the orbits devoted to NICMOS HDF prime proposal 7235. The observations were completed as planned, and no problems were reported. 1.10 Completed NIC/1/2/3 7456 (Water in Protoplanetary Disks) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the protoplanetary disk around stars DR-Tau and Haro-6-37 to search for water in their disks. This survey will be the first to take a detailed look at the gas phase properties of pre-main-sequence disks at AU distances. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Operations Notes: The Battery #2 Capacity Check Test began at 021/1714Z. Discharge began at 02200Z. At 021/204442Z, during an SSR science playlack, error detection and correction counts (EDACs) began accumulating, consistent with the "normal" area of EDAC counts but at a higher rate than usual. The eventual total for that dump activity was 491 counts as opposed to the nominal 120 counts. In a subsequent playback the EDAC counts resumed incrementing. The final count was 517. Later, the entire SSR segment was re-transmitted. The same number of EDAC counts resulted (517). The errors were cleared at 021/224607Z. These events are documented in HSTAR 6325. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH